OpenStudy (anonymous):

What does this equal to y = 2/3 x - 1 y = 3

OpenStudy (anonymous):

are you solving for x?

OpenStudy (anonymous):

mhmmm

OpenStudy (anonymous):

its about Solving Systems of Equations Using Any Method

OpenStudy (anonymous):

if so then \[3 +1 = \frac{ 2 }{ 3 }x\] \[4*3 = \frac{ 2 }{ 3 }x *3 \rightarrow 12 = 2x\] then its clear that x = 6 you can sub that in for x and verify y

OpenStudy (anonymous):

ok ok ty XD

OpenStudy (anonymous):

6x + 7y = 20 y = 2x

OpenStudy (triciaal):

substitute for y in the first equation and solve for x then y = 2x

OpenStudy (triciaal):